# Evolution and rotation of optically bound matter in focused laser beams

### v1

@@ 2.4 Static Electromagnetic Force Simulations: Correlating Torque and Symmetry
Our experimental results demonstrate a clear correlation between the loss of structural symmetry and the cessation of rotation. To computationally validate this finding, we used a generalized multiparticle Mie-theory-based Python package, MiePyparkerOpticalMatterMachines2020, to perform electrodynamic simulations and calculate the optical forces and torques on static assemblies. We focus on the crucial difference in torque generated by two key structural symmetries: the hexagonal packing observed in our stable, rotating assemblies and the square packing that appears as an intermittent, disordered state in our experiments. We can observe that the scattering cross sections of a square and hexagonal array shown in Figure 5a exhibit a clear difference. Thus, trivially, the interaction of these OM arrays with light is significantly different and dependent on the structure.
As established in previous studies, the transfer of torque to an assembly is fundamentally linked to its structural symmetry.chenNegativeOpticalTorque2014, hanCrossoverPositiveNegative2018a, taoRotationalDynamicsIndirect2023a When a photon with an initial angular momentum, $m_{i}$, scatters from a structure possessing a discrete $m_{s}$-fold rotational symmetry, the final angular momentum, $m$, of the scattered photon is constrained to specific allowed channels. The relationship is described by the equation:
@@
Thus, based on our experiments, we computationally investigated the transfer of torque to OM arrays under conditions of a circularly polarized, focused beam. See methods and Supplementary Information S4 for further simulation details. We chose square symmetry for comparison, since many of the intermittent states in the fluid-like state of the assembly have square packing. This is visible in Supplementary Videos 3 and 4, as well as in the snapshots of assemblies in Figures 4 and 2a.
@@
We calculate the orbital torques from these forces and study these torques as the static OM assembly evolves from a single particle to multiple particles. We show in Figures 5d and e the total orbital torque as heatmaps on OM assemblies for different numbers of particles with hexagonal and square packing symmetry, respectively. When a 1064 nm laser is used to create optical matter, the optical binding distance is nearly $\frac{1064}{1.33}\approx 800$ nm. However, since the particles are subject to fluctuations in a fluid, we look at the torque for different lattice spacings. We also studied the torque for these two cases to change the beam waist of the focused beam. The primary effect of increasing the beam waist is that the torque magnitude reduces. The negative torque regime is also shifted to higher interparticle separations. For the hexagonal packing with a lattice spacing of 800 nm, the torque values increase positively with the number of particles. Meanwhile, for square lattice spacings, it is near zero and even negative. For a beam waist of 1 $\mu m$, the heat map shows a band of low and negative torque for four particles as a function of interparticle separation. This shows a strong dependence of the torque transferred on the structure of the OM. For the square-packed assembly, the torque values are predominantly negative and small for lattice spacing of 800 and 1000 nm. For a lattice spacing of 600 nm, the torque is positive but smaller by an order of magnitude compared to the hexagonal assembly. The torque study on static assemblies establishes a dependence of the torque on the structure. Particularly, a sufficient positive torque is only generated in triangular assemblies with a lattice spacing of 800 nm, which is the theoretical optical binding distance given by $d_{OB}=\lambda_{m}=1064/1.33$. And a negative or small positive torque for square assembly with the same parameters. We performed similar simulations to study the behaviour of assemblies formed using different laser beams of gold nanoparticles of different particle sizes. The results are presented in Supplementary Information S5.
@@
Figure 6: Electrodynamic-Langevin dynamic simulations. a) Initial positions for 3P and 4P assembly and the tracks generated by the dynamic simulation. The 3P assembly rotates at a faster rate than the 4P assembly. Additionally, the 4P assembly initially arranged as a square degenerates to a parallelogram with triangular symmetry. The simulations are shown in supplementary video 3. b) Initial positions for 7P and 8P assemblies and the tracks generated by the dynamic simulations. The eight particles are arranged in two ideal representative positions, viz., in a circle and a square arrangement. The 7P arrangement rotates stably. One particle track is marked in purple for visual clarity. The circularly arranged 8P assembly doesn’t rotate around the beam. The square 8P array degenerates and rotates slowly. The simulations are shown in supplementary video 4. c) The 7P dynamic simulation with varying beam widths. The reduced intensity leads to a lower angular velocity. d) The effect of laser power is studied for 7P rotation for three different beam widths. e) The stable optical binding distance is shown for differing beam widths as a function of power.
@@ 2.5 Electrodynamic-Langevin dynamic simulations show the dynamical behaviour
The simulations for the three- and four-particle assemblies, shown in Figure 6a and Supplementary Video 5, provide a dynamic confirmation of our experimental results. The Electrodynamic-Langevin dynamic (EDLD) simulation comparing the triangular 3P assembly and square 4P assembly is shown in Supplementary Video 5. The tracks generated from the simulation are summarised in Figure 6a. The initial positions for both cases are shown on the left. The tracks generated by the time evolution are shown on the right with initial positions marked with blue dots. Both tracks are for the same duration. The 3P assembly rotates stably, maintaining the triangular formation, but the 4P assembly, which started as a square, doesn’t rotate as much at the same time. The square configuration exhibits instability, and in other simulations, it evolves into a triangular arrangement due to stochastic noise. This triangular arrangement shows rotation thereafter.
@@
We extended these dynamic simulations to larger assemblies by modeling the behavior of 7P and 8P arrays. We set up the initial condition for the 7P simulation by placing one particle at the origin and symmetrically placing the six other particles around in a circle. This is shown schematically in the top panel of Figure 6b. For the 8P assembly, we take two separate initial positions similar to the states observed in the experiment (see Figure 4b). Firstly, we consider a circular assembly with one particle in the centre and seven particles symmetrically around a circle of radius 0.8 $\mu m$. Next, we make a square-packed state of 8 particles. The comparison of the evolution for these three states is shown in Supplementary Video 6. The corresponding trajectories are shown in the bottom panel of Figure 6b. The 7P assembly exhibits stable rotation, and the trajectories overlap. The yellow trajectory is overlaid on the other trajectories for clear visualization. The 8P circular assembly doesn’t show orbital rotation as seen for the 7P assembly. In contrast, this assembly slowly orbits the beam axis rigidly, as shown in the corresponding track. In comparison, the square-packed 8P assembly shows instability and rotates slowly after reordering to a hexagonal state. We conclude from the EDLD simulations that triangular or hexagonal states are stable and efficiently harness angular momentum from light in a positive manner. In contrast, the square-packed states are unstable and degenerate into other non-square-packed states, and do not show rotation.
@@
We focus on the stable rotating state of 7 particles and study its properties as the laser power and beam waists are changed in EDLD simulations. In Figure 6c, we show the dependence of the angular velocity of the assembly. The angular velocity increases linearly with power, and changing the beam waist changes the slope of the linear increase. This is expected as increasing the beam waist reduces the intensity quadratically ($I\sim w_{0}^{-2}$) in the sample plane. We also observe that the stable orbiting radius of the surrounding particles varies as a function of the beam waist. We show this in Figure 6d. Generally, reducing the beam waist decreases the stable orbiting radii. This can be explained by the fact that the beam exerts a smaller scattering force on the peripheral particles when the beam waist is smaller. We performed EDLD simulations to investigate the behavior of assemblies composed of particles of varying sizes and materials. The results are presented in Supplementary Information S6.
@@
To conclude, we demonstrated the dynamic evolution of OM assemblies using EDLD simulations. The simulations qualitatively support our experimental results, establishing a connection between the structure of the assembly and the torque transferred to it. While our simulations support the experimental findings, it’s important to discuss key discrepancies that provide valuable insights and avenues for future research. Our dynamic simulations of an 8-particle assembly, starting from a hexagonally packed state, show stable rotation. However, our experiments with a focused 1064 nm laser beam rarely produced a stable 8-particle hexagonal assembly. Instead, the assembly typically exhibited a fluid-like behavior. This suggests that the experimental conditions introduce factors not fully captured by our simplified model. The discrepancy could arise from several unaccounted factors. Focused beams have a strong intensity gradient not just in the radial direction but also along the beam axis. As the assembly grows in size and its size approaches the beam waist, the outermost particles may experience axial forces leading to their out-of-plane fluctuations. These out-of-plane movements could be disrupting the optical binding and leading to the observed ‘fluid-like’ state. Our simulations, which solve the motion in two dimensions only, would not capture this behaviour. Another potential factor we have largely ignored in calculating the force and motion is the presence of the glass cover slip and associated surface effects. We do so, as it has been previously shown that the presence of the surface does not affect the optical forces in similar experiments. Finally, the model solves the dynamical problem with perfectly identical spheres. This assumption may not hold true in the laboratory, as nanoparticle samples often exhibit considerable polydispersity, which could affect the precise arrangement and stability of the optical matter assemblies. Despite these model simplifications, the simulations have provided crucial insights by establishing a clear link between structural symmetry and the transfer of torque in focused beams.
3 Conclusion
@@
We experimentally find that the structure of the assemblies can transition from a hexagonal order to non-hexagonal, fluid-like states as a function of particle and incident beam characteristics. We demonstrate that the assemblies remain stable up to a certain number of particles, depending on the beam width. After a critical size is reached, the addition of another particle, which acts like a defect, leads to geometrically frustrated configurations, creating a chaotic potential energy landscape that lacks a clear rotational pathway. The resulting assembly exists as a kinetically arrested transient phase, fluctuating between multiple low-energy states where the rotational drive is overcome by the loss of symmetry. This study builds on the previously explored OM rotation under broad-beam illumination and also establishes that the behavior of optical matter is fundamentally different and more complex under the strong intensity gradients of a focused beam. We corroborate these experimental results with force calculations on static assemblies showing torque behaviour for assemblies of different symmetries. In particular, we compare triangular and hexagonal packed states for three and seven particles with square packed states for four and eight particles, as well as circularly packed states for eight particles. We also perform electrodynamic-Langevin dynamic simulations, which show that the triangular and hexagonal states rotate stably, while square-packed states are unstable and reorder to a hexagonal state. And the circularly packed eight-particle arrangement doesn’t show any orbital rotation, despite being electrodynamically stable for short times.
@@ 4.3 Particle Tracking
The images captured were tracked using the Trackmate plugin of the FiJi distribution of ImageJschindelinFijiOpenSourcePlatform2012. The images were taken using brightfield illumination, and the gold nanoparticles are visualised as dark objects on a bright background. We invert these images for tracking using the TrackMate plugin. The images extracted from the tracked processed video show particles as white objects on a black background.
@@ 4.4 Simulations
We use a Python-based generalised multiparticle Mie theory package, MiePyparkerOpticalMatterMachines2020, to calculate all the optical forces and scattering cross-sections. We solve the overdamped Langevin equation for various numbers of particles and initial positions in Python. At each time step, the optical force on each particle is calculated using MiePy. The time step for the simulations is 1 $\mu s$.

### v2

@@ 2.4 Static Electromagnetic Force Simulations: Correlating Torque and Symmetry
Our experimental results demonstrate a clear correlation between the loss of structural symmetry and the cessation of rotation. To computationally validate this finding, we used a generalized multiparticle Mie-theory-based Python package, MiePyparker_optical_2020, to perform electrodynamic simulations and calculate the optical forces and torques on static assemblies. We focus on the crucial difference in torque generated by two key structural symmetries: the hexagonal packing observed in our stable, rotating assemblies and the square packing that appears as an intermittent, disordered state in our experiments. We can observe that the scattering cross sections of a square and hexagonal array shown in Figure 5a exhibit a clear difference. Thus, trivially, the interaction of these OM arrays with light is significantly different and dependent on the structure.
As established in previous studies, the transfer of torque to an assembly is fundamentally linked to its structural symmetry.chen_negative_2014, han_crossover_2018, tao_rotational_2023 When a photon with an initial angular momentum, $m_{i}$, scatters from a structure possessing a discrete $m_{s}$-fold rotational symmetry, the final angular momentum, $m$, of the scattered photon is constrained to specific allowed channels. The relationship is described by the equation:
@@
Thus, based on our experiments, we computationally investigated the transfer of torque to OM arrays under conditions of a circularly polarized, focused beam. See methods and Supplementary Information S4supple for further simulation details. We chose square symmetry for comparison, since many of the intermittent states in the fluid-like state of the assembly have square packing. This is visible in Supplementary Videos 3 and 4, as well as in the snapshots of assemblies in Figures 4 and 2a.
@@
We calculate the orbital torques from these forces and study these torques as the static OM assembly evolves from a single particle to multiple particles. We show in Figures 5d and e the total orbital torque as heatmaps on OM assemblies for different numbers of particles with hexagonal and square packing symmetry, respectively. When a 1064 nm laser is used to create optical matter, the optical binding distance is nearly $\frac{1064}{1.33}\approx 800$ nm. However, since the particles are subject to fluctuations in a fluid, we look at the torque for different lattice spacings. We also studied the torque for these two cases to change the beam waist of the focused beam. The primary effect of increasing the beam waist is that the torque magnitude reduces. The negative torque regime is also shifted to higher interparticle separations. For the hexagonal packing with a lattice spacing of 800 nm, the torque values increase positively with the number of particles. Meanwhile, for square lattice spacings, it is near zero and even negative. For a beam waist of 1 $\mu m$, the heat map shows a band of low and negative torque for four particles as a function of interparticle separation. This shows a strong dependence of the torque transferred on the structure of the OM. For the square-packed assembly, the torque values are predominantly negative and small for lattice spacing of 800 and 1000 nm. For a lattice spacing of 600 nm, the torque is positive but smaller by an order of magnitude compared to the hexagonal assembly. The torque study on static assemblies establishes a dependence of the torque on the structure. Particularly, a sufficient positive torque is only generated in triangular assemblies with a lattice spacing of 800 nm, which is the theoretical optical binding distance given by $d_{OB}=\lambda_{m}=1064/1.33$. And a negative or small positive torque for square assembly with the same parameters. We performed similar simulations to study the behaviour of assemblies formed using different laser beams of gold nanoparticles of different particle sizes. The results are presented in Supplementary Information S5.supple
@@
Figure 6: Electrodynamic-Langevin dynamic simulations. a) Initial positions for 3P and 4P assembly and the tracks generated by the dynamic simulation. The 3P assembly rotates at a faster rate than the 4P assembly. Additionally, the 4P assembly initially arranged as a square degenerates to a parallelogram with triangular symmetry. The simulations are shown in supplementary video 3. b) Initial positions for 7P and 8P assemblies and the tracks generated by the dynamic simulations. The eight particles are arranged in two ideal representative positions, viz., in a circle and a square arrangement. The 7P arrangement rotates stably. One particle track is marked in purple for visual clarity. The circularly arranged 8P assembly does not rotate around the beam. The square 8P array degenerates and rotates slowly. The simulations are shown in supplementary video 4. c) The 7P dynamic simulation with varying beam widths. The reduced intensity leads to a lower angular velocity. d) The stable optical binding distance is shown for differing beam widths as a function of power.
@@ 2.5 Electrodynamic-Langevin dynamic simulations show the dynamical behaviour
The simulations for the three- and four-particle assemblies, shown in Figure 6a and Supplementary Video 5, provide a dynamic confirmation of our experimental results. The Electrodynamic-Langevin dynamic (EDLD) simulation comparing the triangular 3P assembly and square 4P assembly is shown in Supplementary Video 5. The tracks generated from the simulation are summarised in Figure 6a. The initial positions for both cases are shown on the left. The tracks generated by the time evolution are shown on the right with initial positions marked with blue dots. Both tracks are for the same duration. The 3P assembly rotates stably, maintaining the triangular formation, but the 4P assembly, which started as a square, does not rotate as much at the same time. The square configuration exhibits instability, and in other simulations, it evolves into a triangular arrangement due to stochastic noise. This triangular arrangement shows rotation thereafter.
@@
We extended these dynamic simulations to larger assemblies by modeling the behavior of 7P and 8P arrays. We set up the initial condition for the 7P simulation by placing one particle at the origin and symmetrically placing the six other particles around in a circle. This is shown schematically in the top panel of Figure 6b. For the 8P assembly, we take two separate initial positions similar to the states observed in the experiment (see Figure 4b). Firstly, we consider a circular assembly with one particle in the centre and seven particles symmetrically around a circle of radius 0.8 $\mu m$. Next, we make a square-packed state of 8 particles. The comparison of the evolution for these three states is shown in Supplementary Video 6. The corresponding trajectories are shown in the bottom panel of Figure 6b. The 7P assembly exhibits stable rotation, and the trajectories overlap. The yellow trajectory is overlaid on the other trajectories for clear visualization. The 8P circular assembly does not show orbital rotation as seen for the 7P assembly. In contrast, this assembly slowly orbits the beam axis rigidly, as shown in the corresponding track. In comparison, the square-packed 8P assembly shows instability and rotates slowly after reordering to a hexagonal state. We conclude from the EDLD simulations that triangular or hexagonal states are stable and efficiently harness angular momentum from light in a positive manner. In contrast, the square-packed states are unstable and degenerate into other non-square-packed states, and do not show rotation.
@@
We focus on the stable rotating state of 7 particles and study its properties as the laser power and beam waists are changed in EDLD simulations. In Figure 6c, we show the dependence of the angular velocity of the assembly. The angular velocity increases linearly with power, and changing the beam waist changes the slope of the linear increase. This is expected as increasing the beam waist reduces the intensity quadratically ($I\sim w_{0}^{-2}$) in the sample plane. We also observe that the stable orbiting radius of the surrounding particles varies as a function of the beam waist. We show this in Figure 6d. Generally, reducing the beam waist decreases the stable orbiting radii. This can be explained by the fact that the beam exerts a smaller scattering force on the peripheral particles when the beam waist is smaller. We performed EDLD simulations to investigate the behavior of assemblies composed of particles of varying sizes and materials. The results are presented in Supplementary Information S6.supple
@@
In this section, we demonstrated the dynamic evolution of OM assemblies using EDLD simulations. The simulations qualitatively support our experimental results, establishing a connection between the structure of the assembly and the torque transferred to it. While our simulations support the experimental findings, it’s important to discuss key discrepancies that provide valuable insights and avenues for future research. Our dynamic simulations of an 8-particle assembly, starting from a hexagonally packed state, show stable rotation. However, our experiments with a focused 1064 nm laser beam rarely produced a stable 8-particle hexagonal assembly. Instead, the assembly typically exhibited a fluid-like behavior. This suggests that the experimental conditions introduce factors not fully captured by our simplified model. The discrepancy could arise from several unaccounted factors. Focused beams have a strong intensity gradient not just in the radial direction but also along the beam axis. As the assembly grows in size and its size approaches the beam waist, the outermost particles may experience axial forces leading to their out-of-plane fluctuations. These out-of-plane movements could be disrupting the optical binding and leading to the observed ‘fluid-like’ state. Our simulations, which solve the motion in two dimensions only, would not capture this behaviour. Another potential factor we have largely ignored in calculating the force and motion is the presence of the glass cover slip and associated surface effects. We do so, as it has been previously shown that the presence of the surface does not affect the optical forces in similar experiments. Finally, the model solves the dynamical problem with perfectly identical spheres. This assumption may not hold true in the laboratory, as nanoparticle samples often exhibit considerable polydispersity, which could affect the precise arrangement and stability of the optical matter assemblies. Despite these model simplifications, the simulations have provided crucial insights by establishing a clear link between structural symmetry and the transfer of torque in focused beams.
3 Discussion and Conclusion
@@
We experimentally find that the structure of the assemblies can transition from a hexagonal order to non-hexagonal, fluid-like states as a function of particle and incident beam characteristics. We demonstrate that the assemblies remain stable up to a certain number of particles, depending on the beam width. After a critical size is reached, the addition of another particle, which acts like a defect, leads to geometrically frustrated configurations, creating a chaotic potential energy landscape that lacks a clear rotational pathway. The resulting assembly exists as a kinetically arrested transient phase, fluctuating between multiple low-energy states where the rotational drive is overcome by the loss of symmetry. This study builds on the previously explored OM rotation under broad-beam illumination and also establishes that the behavior of optical matter is fundamentally different and more complex under the strong intensity gradients of a focused beam. We corroborate these experimental results with force calculations on static assemblies showing torque behaviour for assemblies of different symmetries. In particular, we compare triangular and hexagonal packed states for three and seven particles with square packed states for four and eight particles, as well as circularly packed states for eight particles. We also perform electrodynamic-Langevin dynamic simulations, which show that the triangular and hexagonal states rotate stably, while square-packed states are unstable and reorder to a hexagonal state. And the circularly packed eight-particle arrangement does not show any orbital rotation, despite being electrodynamically stable for short times.
@@ 4.3 Particle Tracking
The images captured were tracked using the Trackmate plugin of the FiJi distribution of ImageJschindelin_fiji_2012. The images were taken using brightfield illumination, and the gold nanoparticles are visualised as dark objects on a bright background. We invert these images for tracking using the TrackMate plugin. The images extracted from the tracked processed video show particles as white objects on a black background.
@@ 4.4 Simulations
We use a Python-based generalised multiparticle Mie theory package, MiePyparker_optical_2020, to calculate all the optical forces and scattering cross-sections. We solve the overdamped Langevin equation for various numbers of particles and initial positions in Python. At each time step, the optical force on each particle is calculated using MiePy. The time step for the simulations is 1 $\mu s$.